OpenStudy (love_to_love_you):

Write each polynomial function in standard form, classify it by degree, and determine the end behavior of its graph. Show all work. 1. y= 12-x^4 2. y= 2x^3-6x+3x^2-x^4+12 3. y= 10-3x^3+3x^2+x^4

OpenStudy (love_to_love_you):

Will fan and medal if you show all steps

OpenStudy (danjs):

What do you know how to do? What is the standard form for a polynomial?

OpenStudy (love_to_love_you):

This is a new thing. I've never done it before so I just need to see the steps to understand. Explaining is harder for me, I know, weird but it's just the way my brain works.

OpenStudy (anonymous):

\[y= 12-x^4\] i would say standard form is \[-x^4+12\]

OpenStudy (anonymous):

degree is 4

OpenStudy (anonymous):

since the leading coefficient is negative the "end behaviour" is \[\lim_{x\to \infty}=\lim_{x\to -\infty}-x^4+12=-\infty\]